Directions: Answer each question to the best of your ability. Please show all work (including the 7 steps of Hypothesis Test) and round decimals to the nearest ten-thousandths5. The authors of the article Accommodating Persons with AIDS: Acceptance and Rejection in Rental Situations (Journal of Ap
Answer #1

Answer:

Sample size, n1 -80 1 Sample proportion, Pl -x1/n1 61/80 0.7625 Sample size, n2 =80 Sample proportion, P2 -x2/n2 =32/80 0.40Test statistic, Z -(p2 -- pl)/SE -(0.40 - 0.7625)/0.078006 -4.65 Critical value(Zcritical): At 0.01 significance level (a), f